OpenStudy (anonymous):

Find an equation of the line tangent to the graph of y=-2/x^2 at the point (2, -2/4)

OpenStudy (campbell_st):

can you find the 1st derivative of the curve..?

OpenStudy (anonymous):

I understand the derivative of the curve is 4/X^3 I just don't know where to go from there.

OpenStudy (campbell_st):

ok... so the 1st derivative gives you the equation for the slope of the tangent at any point on the curve. so substitute x = 2 into the equation to find the slope of the tangent at your point. when you have the slope substitute it into the slope intercept form for the tangent and also substitute x = 2 and y = - 1/2 to find the value of b y = mx + b hope this helps

OpenStudy (anonymous):

Thank you I was making a very stupid mistake. I was going right for the value of b forgetting to plug in 2 for m first. Thanks again

OpenStudy (campbell_st):

glad to help